OpenStudy (anonymous):

Solve for x: the quantity of x plus 4 over 3 = 2

OpenStudy (anonymous):

or try \[\large \frac{x+4}{3}=2\iff x+4=6\]

OpenStudy (anonymous):

subtract \(4\) and get \(x=2\)

OpenStudy (anonymous):

thank you sooo much (: